OpenStudy (anonymous):

How do you write the equation of a line that passes through (2,3) and is perpendicular to y=2x+1

OpenStudy (nikato):

do you know the relation between a line and its perepndicular?

OpenStudy (anonymous):

no i do not

OpenStudy (anonymous):

A perpendicular slope is the negative reciprocal of the slope of the original equation

OpenStudy (anonymous):

Do you know point slope form

OpenStudy (nikato):

well its simple. say you have a line with a slope of 5. its perpendicular buddy would be the negative recipricol which is -1/5

OpenStudy (anonymous):

My teacher doesnt teach me well i just need instructions on how to do it and ill be golden

OpenStudy (nikato):

so what would be the slope of its perpendicular(your equation)

OpenStudy (anonymous):

im not sure honestly

OpenStudy (anonymous):

so the equation for point slope form is \[y - y _{1} = m(x - x _{1})\]m = slope

OpenStudy (nikato):

ok, so do you understand the equation y=mx + b. and thats the equation you were givien, or just in that format